# Distributionally robust monopoly pricing: Switching from low to high prices in volatile markets

### v1

@@ Abstract
Traditional monopoly pricing assumes sellers have full information about consumer valuations. We consider monopoly pricing under limited information, where a seller only knows the mean, variance and support of the valuation distribution. The objective is to maximize expected revenue by selecting the optimal fixed price.
We adopt a distributionally robust framework, where the seller considers all valuation distributions that comply with the limited information. We formulate a maximin problem which seeks to maximize expected revenue for the worst-case valuation distribution. The minimization problem that identifies the worst-case valuation distribution is solved using primal-dual methods, and in turn leads to an explicitly solvable maximization problem. This yields a closed-form optimal pricing policy and a new fundamental principle prescribing when to use low and high robust prices.
We show that the optimal policy switches from low to high prices when variance becomes sufficiently large, yielding significant performance gains compared with existing robust prices that generally decay with market uncertainty. This presents guidelines for when the seller should switch from targeting mass markets to niche markets. Similar guidelines are obtained for delay-prone services with rational utility-maximizing customers, underlining the universality and wide applicability of the low-high pricing principle.
@@ 1 Introduction
Optimizing revenue in contemporary markets represents a pivotal challenge, extensively explored within theoretical pricing models spanning disciplines such as operations research, computer science, and economics. Prevailing in the theoretical pricing literature is the assumption that sellers possess complete knowledge of how the market responds to price variations. This assumption relies on a full understanding of the demand function, serving as a prior in Bayesian analysis, and significantly influencing the derived optimal price. However, the theoretical advantage of assuming full knowledge is compromised by the scarcity or unreliability of market data. A principle known as Wilson’s doctrine Wilson (1987) calls for less detailed yet parsimonious models, and the development of more resilient pricing theories, capable of accommodating imprecise market details Carroll (2019).
Robust pricing theories seek strategies that perform well across diverse markets, reducing sensitivity to detailed market knowledge. The interplay between a seller’s pricing strategy and market knowledge then becomes crucial. A complete lack of knowledge may result in an overly pessimistic worst-case scenario, rendering a robust price excessively conservative and impractical. A more realistic approach is to assume partial knowledge. This paper addresses this scenario in the context of a monopolist selling a single type of good to a diverse, independent consumer base.
The monopolist aims to determine the revenue-maximizing price by balancing aggressive (low) pricing for a larger market share against higher pricing for a smaller market share. Each consumer’s valuation of the good is an independent draw from a given valuation distribution with the purchase occurring when this valuation exceeds the price. The expected revenue is defined as the product of the price times the probability that the valuation exceeds the price. While the optimal price is easily determined when the valuation distribution is known Riley and Zeckhauser (1983), Myerson (1981), the challenge arises when only partial market knowledge is available.
@@
This study focuses on the case where only the mean, variance, and an upper limit of the valuation distribution are known. Mean-variance information proves suitable for robust decision-making, capturing both the anticipated future scenario (mean) and the risk associated with less likely scenarios (variance). This choice is motivated by the often scarce or non-existent data faced by sellers, where mean, variance, and upper bounds can be based on expert opinions—intuitive statistics that are easier to estimate than the entire valuation distribution; see e.g. Scarf (1958), Azar and Micali (2012).
@@
The partially-informed seller needs to decide on the price in the face of uncertain market demand, much in the same way as an investor who only knows the risk and return of assets. To do so, the seller employs a maximin decision framework. The maximin price, or robust price, is determined by first establishing the worst-case valuation distribution that minimizes expected revenue for all prices and then maximizing expected revenue by setting the optimal price. This involves solving two optimization problems—a semi-infinite linear program and finding the global optimum of a bimodal function that depends delicately on the price and information parameters. The comprehensive nature of this approach reveals the intricate interplay between partial knowledge and robust pricing, surpassing earlier studies Azar and Micali (2012), Kos and Messner (2015), Suzdaltsev (2018), Chen et al. (2022a, 2023b) and highlighting a novel robust pricing principle emerging only when knowledge includes both the upper bound and variance.
@@
Before we explain this general principle, we consider classical monopoly pricing with a fully informed monopolist who knows that valuations follow some given cumulative distribution function $F$ with probability density function $f$. For any given price $p$, the expected revenue is $p(1-F(p))$ and the monopolist thus solves the maximization problem
$\sup_{p}p\left(1-F(p)\right).$ (1)
First-order conditions then prescribe that the optimal price $p^{*}$ solves
@@
The “U-shaped” nature (or quasi-convexity) of revenue as function of demand dispersion – as displayed in the right panel of Figure LABEL:fig:beta – is a known phenomenon in monopolistic markets Johnson and Myatt (2006), Sun (2012). Rigorous underpinning is provided in Johnson and Myatt (2006) by an explanatory analysis that shows the U-shape occurs naturally by imposing only mild conditions on the valuation distribution. Consequently, when there is sufficiently low dispersion, indicative of relatively uniform consumer valuations, sellers are inclined to cater to “mass markets” through competitive pricing strategies. Conversely, as dispersion increases, the average consumer’s willingness to pay diminishes. The heightened heterogeneity among consumers prompts firms to adopt a strategy of restricting sales to a more specialized “niche market” characterized by higher prices. We add a theory aligned with the above intuition, showing that for a rich class of potential consumer valuations, a robust seller indeed switches from low pricing with large market share to high pricing and small market share in response to increased consumer valuation dispersion.
@@
To understand the main pricing principle in this paper, now assume that the seller only knows the mean, standard deviation and support, and chooses to fit to these summary statistics a Beta distribution. Since the monopolist now only pretends to know the valuation distribution, the optimal price found from assuming a Beta distribution is only an approximation, as the ground truth valuation is likely not to be Beta distributed. In fact, the Beta distribution is only one of infinitely many distributions with the same mean, standard deviation and support, and only serves to calculate the expected revenue. This paper presents an alternative pricing strategy, where the monopolist does not commit to a specific valuation distribution, and instead first determines the worst-case valuation distribution that minimizes demand, and only then solves for the revenue-maximizing price. We prove that this robust maximin price corresponds to the black solid line in Figure LABEL:fig:beta. As this paper’s title says, the monopolist switches from low to high prices when the variance exceeds a certain threshold. Before that threshold, the maximin price decreases with variance, at the threshold the price jumps to a higher value, and continues to increase after the threshold.
@@
Contributions. We reveal a new phenomenon of abruptly switching from low to high pricing when the market volatility becomes sufficiently large. In this way, we contribute to advancing our understanding of posted pricing in the presence of limited market information. Azar and Micali (2012) pioneered this direction for knowledge of the mean and standard deviation, demonstrating that the optimal posted price maximizing expected revenue in the worst case can be expressed as an explicit function that decays with the standard deviation. Our more general results align with theirs, and their pricing function appears prominently in several of our theorems for low-risk markets. Another relevant study by Kos and Messner (2015) considers information about the mean and bound on support, so the seller lacks knowledge about the variance. Kos and Messner (2015) also derive an explicit pricing function which appears in our study when we consider high-risk markets. We summarize the existing works and our new results in Table 1.
@@
Mean Variance Support Suzdaltsev (2018) This paper: -Proposition 4.1 -Proposition 4.2 -Theorem 5.2 This paper and unpublished work of Suzdaltsev present robust prices that first decrease and then increase with variance, causing the switch from low to high pricing strategies in response to increased market volatility. The high pricing strategy targets niche markets.
@@
From a more theoretical perspective, we rigorously establish several propositions and theorems (see Table 1) for both the classical pricing problem and pricing in an extended setting with delay-prone services. These theoretical results together provide a firm underpinning of this principle of low and high robust pricing. In all cases, we derive the optimal posted prices for when the seller knows the mean, bound on support, and standard deviation range of the valuation distributions, generalizing all earlier works including Azar and Micali (2012) and Kos and Messner (2015). We obtain explicit expressions for the optimal prices by solving a maximin problem with distributionally robust optimization methods. We solve the minimization problem with a classical primal-dual method, see e.g. Rogosinski (1958), Shapiro (2001), and show that the solution renders a bimodal revenue function with up to three candidate prices that achieve local maxima. Solving the maximization problem then requires showing that the global optimum is always one of these three local optima, ruling out boundary points; the robust seller then always chooses one of these three candidate prices, a crucial insight that generalizes all existing contributions listed in Table 1.
Outline. The paper is structured as follows. Section 2 presents an overview of related literature on monopoly pricing and distributionally robust optimization. Section 3 then presents tight tail bounds for the valuation distribution, and Section 4 the first results on robust pricing and the low-high pricing phenomenon. Section 5 contains the most general result for low-high pricing by showing that the partially informed seller always selects one of three closed-form pricing functions. Section 6 treats pricing in delay-prone services and shows that similar robust pricing strategies are optimal. Section 7 concludes and reflects on future work. Appendix A presents the primal-dual proof of the tight-tail bounds and Appendix B contains the proofs of the pricing theorems.
@@ 2 Related literature
Our study adds to the stream of research on robust pricing decisions under limited information, where sellers must commit to fixed prices without possessing complete knowledge of market valuation and the corresponding demand function. Our approach focuses on mean-variance (or standard deviation) information, a widely studied information set in applications such as investment theory Natarajan and Teo (2017), option pricing De Schepper and Heijnen (2007), insurances Roos et al. (2022), service operation management van Eekelen et al. (2022), Taylor (2018) and inventory management Scarf (1958), Gallego (1992). In the management science and operations literature, this information setting is commonly referred to as ‘distribution-free’, highlighting that the decision-making under uncertainty does not require detailed or full distributional knowledge.
@@
Distributionally robust optimization. Such distribution-free analysis combined with optimization falls in the more general domain of distributionally robust optimization (DRO), a mathematical framework that addresses uncertainty in optimization problems by considering an ambiguity set of possible probability distributions for the involved random variables, rather than relying on a single assumed distribution. The ambiguity sets considered in this paper are based on moment information, as in the pioneering work of Scarf (1958) on robust inventory management, and later generalized to a powerful method with many applications Delage and Ye (2010), Wiesemann et al. (2014). Applied to monopoly pricing, this moment-based DRO approach helps create more robust and conservative pricing strategies that perform well under a range of scenarios. Instead of making strong assumptions about the specific valuation distribution, DRO considers a set of possible distributions, forming a certain ambiguity set, and resulting in pricing strategies that are robust across different scenarios, without relying on a single, possibly incorrect, assumed distribution. This aligns well with Occam’s razor Chen et al. (2023a) and Wilson’s doctrine Carroll (2019), two principles of parsimony that aim for explanatory models able to capture the essence without making unnecessary assumptions or requiring overly strong parametric assumptions, such as particular features of probability distributions.
@@
Robust monopoly pricing. Azar and Micali’s foundational result for the information set with mean and variance (and no bound on the support) has been rediscovered and rederived in various papers Carrasco et al. (2018), Chen et al. (2022a), Suzdaltsev (2018). The information set consisting of the triplet mean, variance, and bound on support has remained largely unexplored. A notable exception is Suzdaltsev (2018), where this setting was studied as a special single-bidder case of an auction. In fact, one of the results in this paper (Proposition 4.2) was mentioned without proof in an early draft of Suzdaltsev (2018) and was removed from the final published version. We thus focus on maximizing absolute revenue with information about the first two moments, just as in Azar and Micali (2012), Chen et al. (2022a), Suzdaltsev (2018). Alternative information sets that were studied in the context of robust pricing include knowledge of higher moments Carrasco et al. (2018), percentiles Eren and Maglaras (2010), mean absolute deviation Roos and den Hertog (2021), Elmachtoub et al. (2021), mean preserving contraction and spread Condorelli and Szentes (2020), Roesler and Szentes (2017), or knowing that the valuation distribution is within the proximity of a given reference distribution Bergemann and Schlag (2011), Chen et al. (2023b). Instead of maximin expected revenue, alternative objectives studied in the literature include minimax regret Bergemann and Schlag (2008, 2011), Elmachtoub et al. (2021), Chen et al. (2023b) and the competitive or approximation ratio Eren and Maglaras (2010), Giannakopoulos et al. (2023), Elmachtoub et al. (2021), Wang et al. (2024). Let us also mention that our work connects to Bayesian persuasion Kamenica and Gentzkow (2011), Kamenica (2019), a research line that explores the optimal type or level of information for optimizing individual utility or social welfare.
@@
Regularity assumptions. The lion’s share of the pricing literature assumes strong regularity conditions on the valuation distribution that guarantee existence and uniqueness of an optimal price. In particular, assuming that the valuation has a monotone hazard rate, or some closely related condition, can guarantee that the revenue function is concave and has a unique optimizer Ziya et al. (2004), Schweizer and Szech (2019). Such regularity assumptions greatly simplify the mathematical framework, but true valuation distributions are more often irregular than not Morganti (2022), Roughgarden and Schrijvers (2016). Loertscher and Muir (2022) write: “The standard assumption, which is almost universally maintained in economics, is that expected revenue is concave. The typical justification for this assumption, other than it being standard, is that it is deemed an analytic simplification that permits one to focus on the key economic insights without cluttering the analysis with case distinctions and multiplicity of local maxima. We have never seen it justified on the basis of empirical evidence, and we will not impose it.” See Celis et al. (2014) for empirical evidence that the value distribution need not be regular. See also the literature overview in Loertscher and Muir (2022) for more references on non-concave revenue and a proof that whenever two or more separate markets merge, the revenue is no longer concave. Even in robust settings, regularity conditions are often maintained for mathematical tractability, see e.g. Maskin and Riley (1984), Eren and Maglaras (2010), Allouah et al. (2022). For robust pricing as in this paper, one should let go of regularity completely, as the worst-case distribution that plays a pivotal role in the maximin analysis typically fails this assumption. In fact, if you would add to the minimization problem the condition that the worst-case valuation needs to be regular, we would lose the key take-aways and novel economic insights of this paper.
@@
Pricing in service operations. Monopoly pricing in the context of delay-sensitive service systems has been extensively examined. A well-established body of literature, as detailed by Hassin and Haviv (2003), delves into mathematical models that capture customer decision-making regarding joining delay-sensitive services based on individual utility considerations. The pioneering work by Naor (1969) laid the groundwork for this branch of operations management. Naor’s seminal model views the system as an M/M/1 queue and models customer utility as linearly dependent on the price, value of service, and the anticipated waiting time. Naor revealed that rational individuals have a tendency to overcrowd the system and suggested the introduction of a fixed fee for service that optimizes social welfare. The determination of this socially optimal fee involves solving first-order conditions, akin to classical monopoly pricing. Expanding upon the foundational work by Naor, numerous studies investigated pricing and control considerations for a variety of extended queueing scenarios, with often the focus on determining the price that maximizes revenue or welfare; see e.g. Dewan and Mendelson (1990), Chen and Frank (2004), Ha (1998), Taylor (2018), Baron et al. (2023), Chen et al. (2022b). Stidham (1992) made a significant contribution by introducing the concept of random valuation for individual customers and assuming that the arrival rate of customers with a valuation surpassing a specified threshold. In this framework, valuation is often referred to as willingness-to-pay, and as in classical monopoly pricing, is a random variable with some given distribution. Cachon and Feldman (2011), Haviv and Randhawa (2014) build on Stidham’s framework and assume that a rational customer joins when its random valuation exceeds the sum of the price and the expected delay costs, and determined optimal prices when the valuation distribution is assumed to be known. We will perform a distribution-free analysis of this rational queueing model and show that again the optimal pricing strategy switches from low to high pricing when the standard deviation of the valuation crosses a certain threshold.

### v2

@@ Abstract
Problem definition: Traditional monopoly pricing assumes sellers have full information about consumer valuations. We consider monopoly pricing under limited information, where a seller only knows the mean, variance and support of the valuation distribution. The objective is to maximize expected revenue by selecting the optimal fixed price. Methodology/results: We adopt a distributionally robust framework, where the seller considers all valuation distributions that comply with the limited information. We formulate a maximin problem which seeks to maximize expected revenue for the worst-case valuation distribution. The minimization problem that identifies the worst-case valuation distribution is solved using primal-dual methods, and in turn leads to an explicitly solvable maximization problem. This yields a closed-form optimal pricing policy and a new fundamental principle prescribing when to use low and high robust prices.
Managerial implications: We show that the optimal policy switches from low to high prices when variance becomes sufficiently large, yielding significant performance gains compared with existing robust prices that generally decay with market uncertainty. This presents guidelines for when the seller should switch from targeting mass markets to niche markets. Similar guidelines are obtained for delay-prone services with rational utility-maximizing customers, underlining the universality and wide applicability of the novel pricing policy.
@@ 1 Introduction
Optimizing revenue in contemporary markets represents a pivotal challenge, extensively explored within theoretical pricing models spanning disciplines such as operations research, computer science, and economics. Prevailing in the theoretical pricing literature is the assumption that sellers possess complete knowledge of how the market responds to price variations. This assumption relies on a full understanding of the demand function, serving as a prior in Bayesian analysis, and significantly influencing the derived optimal price. However, the theoretical advantage of assuming full knowledge is compromised by the scarcity or unreliability of market data. A principle known as Wilson’s doctrine calls for less detailed yet parsimonious models, and the development of more resilient pricing theories, capable of accommodating imprecise market details (Wilson 1987, Carroll 2019). Robust pricing theories seek strategies that perform well across diverse markets, reducing sensitivity to detailed market knowledge. The interplay between a seller’s pricing strategy and market knowledge then becomes crucial. A complete lack of knowledge may result in an overly pessimistic worst-case scenario, rendering a robust price excessively conservative and impractical. A more realistic approach is to assume partial knowledge. This paper addresses this scenario in the context of a monopolist selling a single type of good to a diverse, independent consumer base.
@@
The monopolist aims to determine the revenue-maximizing price by balancing aggressive (low) pricing for a larger market share against higher pricing for a smaller market share. Each consumer’s valuation of the good is an independent draw from a given valuation distribution with the purchase occurring when this valuation exceeds the price. The expected revenue is defined as the product of the price times the probability that the valuation exceeds the price. While the optimal price is easily determined when the valuation distribution is known (Riley and Zeckhauser 1983, Myerson 1981), the challenge arises when only partial market knowledge is available. This study focuses on the case when only the mean, variance, and an upper limit of the valuation distribution are known. Mean-variance information proves suitable for robust decision-making, capturing both the anticipated future scenario (mean) and the risk associated with less likely scenarios (variance). This choice is motivated by the often scarce or non-existent data faced by sellers, where the mean, variance, and support upper bound can be based on expert opinions—intuitive statistics that are easier to estimate than the entire valuation distribution; see e.g. Scarf (1958), Azar and Micali (2012).
@@
The partially informed seller needs to decide on the price in the face of uncertain market demand, much in the same way as an investor who only knows the risk and return of assets. To do so, the seller employs a maximin decision framework. The maximin price, or robust price, is determined by first establishing the worst-case valuation distribution that minimizes expected revenue for all prices and then maximizing expected revenue by setting the optimal price. This involves solving two optimization problems—a semi-infinite linear program and finding the global optimum of a bimodal function that depends delicately on the price and information parameters. The comprehensive nature of this approach reveals the intricate interplay between partial knowledge and robust pricing, surpassing earlier studies (Azar and Micali 2012, Kos and Messner 2015, Suzdaltsev 2018, Chen et al. 2022a, 2024) and highlighting a novel robust pricing principle that emerges only when knowledge includes both the support upper bound and variance.
@@
Classical versus robust pricing. Before we explain this general principle, consider classical monopoly pricing with a fully informed monopolist who knows that valuations follow some cumulative distribution function $F$ with probability density function $f$. For any given price $p$, expected revenue is p(1F(p))p(1-F(p)) and the monopolist thus solves the maximization problem $\sup_{p}p\left(1-F(p)\right)$. First-order conditions then prescribe that the optimal price $p^{*}$ solves
@@
The “U-shaped” nature (or quasi-convexity) of revenue as function of demand dispersion – as displayed in the right panel of Figure LABEL:fig:beta – is a known phenomenon in monopolistic markets (Johnson and Myatt 2006, Sun 2012). Rigorous underpinning is provided in Johnson and Myatt (2006) by an explanatory analysis that shows the U-shape occurs naturally by imposing only mild conditions on the valuation distribution. Consequently, when there is sufficiently low dispersion, indicative of relatively uniform consumer valuations, sellers are inclined to cater to “mass markets” through competitive pricing strategies. Conversely, as dispersion increases, the average consumer’s willingness to pay diminishes. The increased heterogeneity among consumers prompts firms to adopt a strategy of restricting sales to a more specialized “niche market characterized by higher prices. We add a theory aligned with the above intuition, showing that for a rich class of potential consumer valuations, a robust seller indeed switches from low pricing with large market share to high pricing and small market share in response to increased consumer valuation dispersion.
@@
To build some first intuition for the robust pricing in this paper, assume that the seller only knows the mean, variance, and support upper bound, and chooses to fit to these summary statistics a beta distribution. Since the monopolist now only pretends to know the valuation distribution, the optimal price found from assuming a beta distribution is only an approximation, as the ground truth valuation is likely not to be beta distributed. In fact, the beta distribution is only one of infinitely many distributions with the same mean, variance and support, and only serves to calculate the expected revenue. This paper presents an alternative pricing strategy, where the monopolist does not commit to a specific valuation distribution, and instead first determines the worst-case valuation distribution that minimizes demand, and only then solves for the revenue-maximizing price. We prove that this robust maximin price corresponds to the black solid line in Figure LABEL:fig:beta. As this paper’s title says, the monopolist switches from low to high prices when the variance exceeds a certain threshold. Before that threshold, the maximin price decreases with variance, at the threshold the price jumps to a higher value, and continues to increase after the threshold.
Worst-case demand implies segmentation. It turns out that the structure of the worst-case valuation distribution identified in this study suggests implicit market segmentation. Specifically, this worst-case market can be interpreted as comprising a low-end segment of customers willing to pay only up to a certain amount and a high-end segment willing to pay significantly more. While detailed knowledge of these segments might lead a seller to consider third-degree price discrimination by setting distinct prices for each group, the seller in this setting only has limited information available. Consequently, price differentiation is not addressed here; the analysis focuses on determining the optimal uniform price for all customers. Due to customers’ perceptions of fairness, however, firms often adopt uniform pricing even in markets with an underlying segmented structure. DellaVigna and Gentzkow (2019) provide substantial evidence that food and mass-merchandise chains often charge nearly uniform prices across stores, despite significant variations in consumer demographics. Beyond fairness perceptions, uniform pricing can perform well in terms of profitability under specific conditions in segmented markets. When the revenue functions for each segment are concave and share a common support (i.e., the same price range applies across all segments), Bergemann et al. (2022) demonstrate that uniform pricing captures at least 50% of the maximum profits achievable under price discrimination. This result holds irrespective of the number of segments or the specific discriminatory prices. However, this profit guarantee collapses when the assumption of common support is violated, underscoring that the degree of market segmentation—particularly differences in price support—has a critical impact on the effectiveness of uniform pricing.
Jumps in sticky prices. Our analysis focuses on a firm selling to a large market without detailed knowledge of the demand function or its potential segmentation. Adopting a robust approach, the firm determines an optimal yet ”sticky” uniform price based on the mean, variance, and support of demand. In the worst-case scenario, demand corresponds to a segmented market where segments differ in price support, making the revenue function fall outside the scope of Bergemann et al. (2022). These differences in support provide strong incentives to adjust the uniform price to better target certain segments. However, such adjustments must be carefully balanced against the risks of customer anger and brand equity concerns when revising prices. For instance, in 2023 and 2024, Tesla implemented significant price reductions across its flagship models, marking a strategic shift from targeting niche markets of early adopters to a broader mass-market audience (Reuters 2024). By lowering entry barriers, Tesla aimed to attract middle-income buyers previously excluded from the electric vehicle market. However, these abrupt price cuts provoked anger among existing Tesla owners, who felt cheated and faced reduced resale values for their vehicles. Despite this backlash, the strategic objective of expanding market reach was prioritized. In our analysis, which solely focuses on revenue maximization and disregards factors such as customer anger and brand reputation, we also identify the triggers for when significant price reductions become profitable, particularly in targeting lower-end segments; in Figure LABEL:fig:beta this trigger is the threshold value for the variance, where the optimal price jumps for low to high or vice versa.
@@
Contributions. We reveal a new phenomenon of abruptly switching from low to high pricing when the market segmentation becomes sufficiently large. In this way, we contribute to advancing our understanding of uniform pricing in the presence of limited market information. Azar and Micali (2012) pioneered this direction for knowledge of the mean and variance, demonstrating that the optimal uniform price maximizing expected revenue in the worst case can be expressed as an explicit function that decays with the variance. Our more general results align with theirs, and their pricing function appears prominently in several of our theorems for low-risk markets. Another relevant study by Kos and Messner (2015) considers information about the mean and bound on support, so the seller lacks knowledge about the variance. Kos and Messner (2015) also derive an explicit pricing function which appears in our study when we consider high-risk markets. We summarize the existing works and our new results in Table 1.
@@
Mean Variance Support Suzdaltsev (2018) This paper: -Theorems 3.1 and 3.2 -Proposition 4.1 and 4.2 -Theorems 5.1, 5.2, 5.3, 5.4 This paper and unpublished work of Suzdaltsev present robust prices that first decrease and then increase with variance, causing the switch from low to high pricing strategies in response to increased market volatility. The high pricing strategy targets niche markets.
@@
From a more theoretical perspective, we rigorously establish several propositions and theorems (see Table 1) for both the classical pricing problem and pricing in an extended setting with delay-prone services. These theoretical results together provide a firm underpinning of this principle of low and high robust pricing. In all cases, we derive the optimal uniform prices for when the seller has full information of the mean, a support upper bound, and a variance range of the valuation distribution, generalizing all earlier works including Azar and Micali (2012), Kos and Messner (2015). We obtain explicit expressions for the optimal prices by solving a maximin problem with distributionally robust optimization methods. We solve the minimization problem with a classical primal-dual method, see e.g. Rogosinski (1958), Shapiro (2001), and show that the solution renders a bimodal revenue function with up to three candidate prices that achieve local maxima. Solving the maximization problem then requires showing that the global optimum is always one of these three local optima, ruling out boundary points; the robust seller then always chooses one of these three candidate prices, a crucial insight that generalizes all existing contributions listed in Table 1.
Outline. The paper is structured as follows. Section 2 presents an overview of related literature on monopoly pricing and distributionally robust optimization. Section 3 obtains the worst-case market in terms of tight tail bounds for the valuation distribution, and presents the main result for low-high pricing by showing that the partially informed seller always selects one of three closed-form pricing functions. This result is proven in Section 4, starting with first presenting results on robust pricing and the low-high pricing phenomenon for special cases. Section 5 treats pricing in delay-prone services and shows that similar robust pricing strategies are optimal. Section 6 concludes and reflects on future work.
@@ 2 Related literature
Our study adds to the stream of research on robust pricing decisions under limited information, where sellers must commit to fixed prices without possessing complete knowledge of market valuation and the corresponding demand function. Our approach focuses on mean-variance information, a widely studied information set in applications such as investment theory (Natarajan and Teo 2017), option pricing (De Schepper and Heijnen 2007), insurances (Roos et al. 2022), service operation management (van Eekelen et al. 2022, Taylor 2018) and inventory management (Scarf 1958, Gallego 1992). In the management science and operations literature, this information setting is commonly referred to as ‘distribution-free’, highlighting that the decision-making under uncertainty does not require detailed or full distributional knowledge.
@@
Distributionally robust optimization. Such distribution-free analysis combined with optimization falls in the more general domain of distributionally robust optimization (DRO), a mathematical framework that addresses uncertainty in optimization problems by considering an ambiguity set of possible probability distributions for the involved random variables, rather than relying on a single assumed distribution. The ambiguity sets considered in this paper are based on moment information, as in the pioneering work of Scarf (1958) on robust inventory management, and later generalized to a powerful method with many applications (Delage and Ye 2010, Wiesemann et al. 2014). Applied to monopoly pricing, this moment-based DRO approach helps create more robust and conservative pricing strategies that perform well under a range of scenarios. Instead of making strong assumptions about the specific valuation distribution, DRO considers a set of possible distributions, forming a certain ambiguity set, and resulting in pricing strategies that are robust across different scenarios, without relying on a single, possibly incorrect, assumed distribution. This aligns well with Occam’s razor (Chen et al. 2023) and Wilson’s doctrine (Carroll 2019), two principles of parsimony that aim for explanatory models able to capture the essence without making unnecessary assumptions or requiring overly strong parametric assumptions.
@@
Robust monopoly pricing. Azar and Micali’s foundational result for the information set with mean and variance (and no bound on the support) has been rediscovered and rederived in various papers (Carrasco et al. 2018, Chen et al. 2022a, Suzdaltsev 2018). The information set consisting of the triplet mean, variance, and bound on support has remained largely unexplored. A notable exception is Suzdaltsev (2018), where this setting was studied as a special single-bidder case of an auction. In fact, one of the results in this paper (Proposition 4.2) was mentioned without proof in an early draft of Suzdaltsev (2018) and was removed from the final published version. We thus focus on maximizing expected revenue with information about the first two moments, just as in Azar and Micali (2012), Chen et al. (2022a), Suzdaltsev (2018). Alternative information sets that were studied in the context of robust pricing include knowledge of higher moments (Carrasco et al. 2018), percentiles (Eren and Maglaras 2010), mean absolute deviation (Roos and den Hertog 2021, Elmachtoub et al. 2021), mean preserving contraction and spread (Condorelli and Szentes 2020, Roesler and Szentes 2017), or knowing that the valuation distribution is within the proximity of a given reference distribution (Bergemann and Schlag 2011, Chen et al. 2024). Instead of maximin expected revenue, alternative objectives studied include minimax regret (Bergemann and Schlag 2008, 2011, Elmachtoub et al. 2021, Chen et al. 2024) and the competitive or approximation ratio (Eren and Maglaras 2010, Giannakopoulos et al. 2023, Elmachtoub et al. 2021, Wang et al. 2024). Let us also mention that our work connects to Bayesian persuasion (Kamenica and Gentzkow 2011, Kamenica 2019), a research line that explores the optimal type or level of information for optimizing individual utility or social welfare.
@@
Regularity assumptions. The lion’s share of the pricing literature assumes strong regularity conditions on the valuation distribution that guarantee existence and uniqueness of an optimal price. In particular, assuming that the valuation has a monotone hazard rate, or some closely related condition, can guarantee that the revenue function is concave and has a unique optimizer (Ziya et al. 2004, Schweizer and Szech 2019). Such regularity assumptions greatly simplify the mathematical framework, but true valuation distributions are more often irregular than not (Morganti 2022, Roughgarden and Schrijvers 2016). Loertscher and Muir (2022) write: “The standard assumption, which is almost universally maintained in economics, is that expected revenue is concave. The typical justification for this assumption, other than it being standard, is that it is deemed an analytic simplification that permits one to focus on the key economic insights without cluttering the analysis with case distinctions and multiplicity of local maxima. We have never seen it justified on the basis of empirical evidence, and we will not impose it.” See Celis et al. (2014) for empirical evidence that the value distribution need not be regular. See also the literature overview in Loertscher and Muir (2022) for more references on non-concave revenue and a proof that whenever two or more separate markets merge, the revenue is no longer concave. Even in robust settings, regularity conditions are often maintained for mathematical tractability, see e.g. Maskin and Riley (1984), Eren and Maglaras (2010), Allouah et al. (2022). For robust pricing as in this paper, one should let go of regularity completely, as the worst-case distribution that plays a pivotal role in the maximin analysis typically fails this assumption. In fact, if you would add to the minimization problem the condition that the worst-case valuation needs to be regular, we would lose the key take-aways and novel economic insights of this paper.
@@
Pricing in service operations. Monopoly pricing in the context of delay-sensitive service systems has been extensively examined. A well-established body of literature, as detailed by Hassin and Haviv (2003), delves into mathematical models that capture customer decision-making regarding joining delay-sensitive services based on individual utility considerations. The pioneering work by Naor (1969) laid the groundwork for this branch of operations management. Naor’s seminal model views the system as an M/M/1 queue and models customer utility as linearly dependent on the price, value of service, and the anticipated waiting time. Naor revealed that rational individuals have a tendency to overcrowd the system and suggested the introduction of a fixed fee for service that optimizes social welfare. The determination of this socially optimal fee involves solving first-order conditions, akin to classical monopoly pricing. Expanding upon the foundational work by Naor, numerous studies investigated pricing and control considerations for a variety of extended queueing scenarios, with often the focus on determining the price that maximizes revenue or welfare; see e.g. Dewan and Mendelson (1990), Chen and Frank (2004), Ha (1998), Taylor (2018), Baron et al. (2023), Chen et al. (2022b). Stidham (1992) made a significant contribution by introducing the concept of random valuation for individual customers and assuming that the arrival rate of customers with a valuation surpassing a specified threshold. In this framework, valuation is often referred to as willingness-to-pay, and as in classical monopoly pricing, is a random variable with some given distribution. Cachon and Feldman (2011), Haviv and Randhawa (2014) build on Stidham’s framework and assume that rational customers join when their random valuation exceeds the sum of the price and the expected delay costs. This requires specifying the valuation distribution in order to determine the optimal price. We will perform a distribution-free analysis of this rational queueing model and show that again the optimal pricing strategy switches from low to high pricing when the variance crosses a certain threshold. Another stream of research on rational queues assumes that customers are heterogeneous in their delay sensitivity rather than valuation (Afèche 2013, Ata and Peng 2018, Yu et al. 2013, Aksin et al. 2013, Nazerzadeh and Randhawa 2018). We will also establish the low-high pricing phenomenon for this setting with a partially known delay-sensitivity distribution.